OpenStudy (anonymous):

A wire 24 inches long is to be cut into four pieces to form a rectangle whose shortest side has a length of x. a. Express the area A of the rectangle as a function of x. b. Determine the domain of of the function

OpenStudy (anonymous):

since the shortest side is x, do you know what is the length of the other side?

OpenStudy (anonymous):

24-x?

OpenStudy (anonymous):

close but it is actually (12-x).... do you know why?

OpenStudy (anonymous):

for a. I got 12-x^2 but I don't know how to get b

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

(shortest side) + (other side) = 12 since the perimeter is 24, then: 2(shortest side) + 2(other side) = 24

OpenStudy (anonymous):

so now, to get the area, just multiply: Area = (shortest side) * (other side) what u got now?

OpenStudy (anonymous):

x(12-x)=12x-x^2

OpenStudy (anonymous):

how do i find the domain?

OpenStudy (anonymous):

good... Area = 12x - x^2

OpenStudy (anonymous):

now, what do you think is the domain?

OpenStudy (anonymous):

here's a hint... what can x be?

OpenStudy (anonymous):

uhhh how would you solve it? would it be all real numbers because i plugged it into my graphing calculator and it looks like it is

OpenStudy (anonymous):

definitely not... x cannot be 30 because the wire is 24 inches long.....

OpenStudy (anonymous):

so x cannot be all real numbers...

OpenStudy (anonymous):

remember what x represents?

OpenStudy (anonymous):

the shorter side of the rectangle

OpenStudy (anonymous):

and how long can that be?

OpenStudy (anonymous):

ohhh so would that mean that its like less than 12

OpenStudy (anonymous):

so would it be like [1,12)

OpenStudy (anonymous):

correct...:) but what?

OpenStudy (anonymous):

oops (0,12) ?

OpenStudy (anonymous):

not [1, 12)

OpenStudy (anonymous):

i agree with the left part of (0, 12)... but not that right part "12)"... do you know why it cant be something like 8?

OpenStudy (anonymous):

OHHHHHH BECAUSE THERES TWO SIDES

OpenStudy (anonymous):

SO ITS LESS THAN SIX?

OpenStudy (anonymous):

very good!!!

OpenStudy (anonymous):

GODDAMMIT RECTANGLES SUCK

OpenStudy (anonymous):

so what's the domain?

OpenStudy (anonymous):

(0,6)

OpenStudy (anonymous):

nicely done... :)

OpenStudy (anonymous):

omg thanks

OpenStudy (anonymous):

yw....:)